OpenStudy (anonymous):

Use fundamental identities to simplify the expression below and then determine which of the following is NOT equivalent cotβsecβ answer choices: 1/sinβ secβ/tanβ 1/(cosβtanβ) secβ cscβ

OpenStudy (anonymous):

I know the answer is not cscβ

OpenStudy (jdoe0001):

well, cot = cos/sin and , sec = 1/cos so multiply them :)

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{cos(x)}}{sin(x)}\cfrac{1}{\cancel{cos(x)}}\)

OpenStudy (anonymous):

If \(\csc\beta\) isn't an answer, then you also know that \(\dfrac{1}{\sin\beta}\) isn't an answer, either.

OpenStudy (anonymous):

secβ/tanβ then?

OpenStudy (anonymous):

Nope,\[\cot\beta\sec\beta=\frac{1}{\tan\beta}{\sec\beta}=\frac{\sec\beta}{\tan\beta}\]

OpenStudy (anonymous):

ok then 1/(cosβtanβ)? because secβ is actually in the problem

OpenStudy (anonymous):

If you rewrite the secant in my last comment, you'll see that they're the same: \[\frac{\sec\beta}{\tan\beta}=\frac{\frac{1}{\cos\beta}}{\tan\beta}=\frac{1}{\cos\beta\tan\beta}\] So you've eliminated four of the answers, meaning the right answer is...

OpenStudy (anonymous):

ha ok secβ, sorry for my stupidity but thanks for help!

OpenStudy (anonymous):

You're welcome!